if you drop a stone into a well, it takes 6 seconds until you hear the splash. How deep is the well?
Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(24713)



Please post your Physics problems on a Physics homework help site.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Alan3354(69443)

if you drop a stone into a well, it takes 6 seconds until you hear the splash. How deep is the well?
---------------------
The speed of sound is very small compared to the time taken to hit the surface of the water. Since you didn't spec a temperature of the ambient air or any equation to use for the time for the sound to travel:
s = -16t^2 = -16*36 = -576 feet
s = -4.9t^2 = -4.9*36 = -176.4 meters